When 0.3860.386 g of sodium metal is added to an excess of hydrochloric acid, 40104010 J of heat are produced. What is the enthalpy of the reaction as written?

2Na(s)+2HCl(aq)⟶2NaCl(aq)+H2(g)2Na(s)+2HCl(aq)⟶2NaCl(aq)+H2(g)
Answer #1

Molar mass of Na = 22.99 g/mol

mass(Na)= 0.386 g

use:

number of mol of Na,

n = mass of Na/molar mass of Na

=(0.386 g)/(22.99 g/mol)

= 1.679*10^-2 mol

For,

1.679*10^-2 mol Na, delta H = -4010 J

In the reaction 2 moles of Na are reacting.

So,

For 2 mol Na, delta H = -4010*2 / (1.679*10^-2) J

= 4.777*10^5 J

= 4.777*10^2 KJ

Answer: 4.777*10^2 KJ